# Optimizing a Worldwide-Scale Shipper Transportation Planning in a Carmaker Inbound Supply Chain

### v2

@@ Abstract
We study the shipper-side design of large-scale inbound transportation networks, motivated by Renault’s global supply chain. We introduce the Shipper Transportation Design Problem, which integrates consolidation, routing, and regularity constraints, and propose a tailored Iterated Local Search (ILS) metaheuristic. The algorithm combines large-neighborhood search with MILP-based perturbations and exploits bundle-specific decompositions and giant container bounds to obtain scalable lower bounds and effective benchmarks. Computational experiments on real industrial data show that the ILS achieves an average gap of 7.9% to the best available lower bound on world-scale instances with more than 700,000 commodities and 1,200,000 arcs, delivering solutions showing a potential of 23.2% cost reduction compared to the Renault-based benchmark. To our knowledge, this is the first approach to solve shipper-side transportation design problems at such scale. Our analysis further yields managerial insights: accurate bin-packing models are essential for realistic consolidation, highly regular plans offer the best balance between cost and operational stability, and outsourcing is only attractive in low-volume contexts, while large-scale networks benefit from in-house planning.
@@ 1 Introduction
Renault exemplifies such a case. As a global car manufacturer with a vast supplier base and worldwide production footprint, Renault moves millions of parts across continents each year. This results in an inbound logistics network of exceptional size and complexity—comprising thousands of suppliers, hundreds of logistics platforms, and multiple industrial sites—requiring highly detailed and scalable optimization methods to ensure efficient operations. In this context, the company cannot rely on off-the-shelf tools or generic outsourcing models. Instead, it requires bespoke, high-performance optimization tailored to its specific needs.
This paper arises from a close industrial collaboration with Renault and addresses a real-world problem of strategic importance: the long-term planning of the company’s inbound transportation network. In this context, our objective is to produce a provisional transportation plan for parts to be assembled in the next six months. Each commodity—defined as a part to be delivered from a specific supplier to a specific industrial site by a given date—must be routed through Renault’s logistics network in a cost-efficient manner. The decision involves choosing whether to ship parts directly or to consolidate them through intermediate platforms.
To allow for optimized decision-making in this context, we formalize the underlying Shipper Transportation Planning Problem (STPP), tailored to the needs of a high-volume industrial shipper, and develop a scalable algorithmic framework, capable of generating high-quality plans on real data from one of Europe’s largest manufacturers.
@@ 1.1 State of the art
The design and optimization of transportation networks have been long-standing challenges in operations research, especially in the context of logistics and supply chain management. The foundational modeling framework is that of multicommodity network flow problems, extensively studied since the 1960s (Ahuja et al., 1988). Early research already recognized the critical importance of flow consolidation in freight transportation (Powell and Sheffi, 1983). More recently, Crainic et al. (2021) provided a comprehensive overview of methods for solving two closely related problems: multicommodity multifacility network design and the load plan design problem. While these models are more tractable, they exhibit many of the same structural and computational challenges as our setting.
@@
Exact solution methods for such problems typically rely on mathematical programming techniques, including decomposition methods such as Lagrangian relaxation, Benders decomposition, and Dantzig-Wolfe reformulation (Crainic and Gendron, 2020). These methods are effective on moderate-size instances, involving a few hundred nodes and commodities, but struggle to scale. Studies by Gendron and Larose (2014) and Frangioni et al. (2017) report computation times of several hours even on such modest networks.
To tackle larger instances, researchers have focused on heuristic and matheuristic methods. Metaheuristics like large neighborhood search and population-based methods have shown promising performance when optimizing realistic-sized networks (Crainic and Gendreau, 2021; Kazemzadeh et al., 2022; Gendron et al., 2018; Paraskevopoulos et al., 2016). For example, in the context of road transportation planning, Bakir et al. (2021), Erera et al. (2013) as well as Lindsey et al. (2016) have addressed networks with tens of thousands of nodes and hundreds of thousands of arcs using sophisticated local descent-based matheuristics. Recently, Eom et al. (2025) proposed a recursive partitioning and batching framework to improve scalability and computation time.
However, these existing approaches face severe limitations when being applied to shipper-side planning at the scale encountered in our context. While some algorithms have been tested on networks of comparable size, they typically address only tens of thousands of commodities, work on time horizons ranging from a day to a week and do not consider explicit bin-packing constraints. This omission is crucial: once bin-packing is included, standard mathematical programming models become intractably large, while decomposition subproblems become too complex or numerous, and existing matheuristics fail to scale. This highlights a major gap in the literature: while LPDPs and large-scale network design problems have been studied from the perspective of carriers and 3PLs, the shipper’s perspective—particularly with explicit consolidation and massive commodity volume—remains underexplored.
@@ 1.2 Contributions
This paper studies the shipper-side design of large-scale inbound transportation networks, focusing on the strategic planning problem faced by Renault. In this context, we contribute (i) a novel, scalable metaheuristic framework that advances the methodological frontier of shipper-side network design, (ii) the first computational results on instances of genuine global industrial scale, and (iii) a set of managerial guidelines that directly inform decision-making in large-scale inbound logistics.
Methodologically, we develop a novel Iterated Local Search (ILS) metaheuristic tailored to the Shipper Transportation Design Problem. The approach integrates efficient large-neighborhood search operators with a perturbation scheme based on tractable MILP relaxations. A key innovation lies in exploiting bundle-specific decompositions and giant container relaxations to obtain scalable lower bounds, which are further used to design a rounding heuristic benchmark. To our knowledge, this is the first approach that combines these methodological ingredients into a unified framework capable of handling networks of industrial scale.
Empirically, we demonstrate that the proposed ILS achieves substantial performance improvements over both Renault’s current planning solutions and a suite of established benchmark heuristics. On the largest “World” instances, the algorithm closes the gap to the best available lower bound to within 7.9% and delivers solutions that reveals a 23.2% potential for Renault’s operational plans cost reduction. Importantly, this represents the first computational study that solves a shipper-side transportation design problem at a scale solving instances with more than 700,000 commodities and 1,200,000 arcs.
Beyond methodological advances, our analysis provides actionable managerial insights. First, we show that a faithful representation of bin-packing consolidation is indispensable for realistic network design, as simplified capacity approximations lead to substantial distortions. Second, we highlight the role of transport regularity: highly regular plans not only reduce computational complexity but also align with planners’ preference for operational stability, while more flexible plans yield only marginal additional cost savings. Third, we analyze outsourcing decisions and reveal a clear volume-dependent threshold: outsourcing may be attractive in small or low-volume networks, but for global, high-volume supply chains such as Renault’s, in-house planning and execution consistently outperform.
@@ 2 Case study & Problem Setting
This section introduces the empirical and operational context underlying the transportation planning problem addressed in this paper. We focus on the inbound supply network of Renault, a global automotive manufacturer operating a high-volume, multi-tier logistics system. The aim is to characterize the network topology, flow structure, and planning requirements that shape the problem formulation developed in the subsequent sections. We begin by describing the global supply network’s physical architecture, highlighting its hub-and-spoke structure and consolidation logic. We then analyze key flow characteristics—volume patterns, shipment granularity, and cost distributions—to identify the core logistical challenges. Finally, we present an informal problem statement that synthesizes these elements into a shipper-side transportation planning problem, capturing the scale, constraints, and cost trade-offs specific to this context.
@@ 2.1 Supply Network
These maps highlight the global scale and complexity of Renault’s supply network, which comprises more than 3,000 sites, 40 industrial sites, and 100 logistics platforms. The term platforms refers to any type of logistics facility where material flows can be consolidated or unconsolidated as needed. The network is predominantly concentrated in Europe, which accounts for 70% of all sites, and consists mainly of suppliers, which account for 90% of all sites in the network. Operating at this scale introduces substantial logistical challenges. Lead times can vary from days to several weeks, depending on the route and transport mode. Intercontinental coordination, sensitivity to global disruptions (e.g., port delays, geopolitical events), and the need to synchronize flow consolidation across multiple tiers (a long-haul core network of platforms and a vast periphery of suppliers and industrial sites), all add to the complexity of transportation planning.
The network exhibits a hub-and-spoke structure: suppliers form a sparsely connected periphery, each typically linked to only one or two other nodes, while platforms act as high-degree hubs. These platforms enable flow consolidation and serve as key transshipment points that reduce the number of required direct supplier–site connections. As a result, the core network of platforms and industrial sites, although relatively small in size, is densely connected and serves as the primary backbone for flow consolidation and routing. This hub-and-spoke structure is a common design principle in large-scale industrial supply networks. By concentrating inbound flows through logistics platforms, firms can exploit economies of scale in transportation, reduce the number of direct connections between suppliers and industrial sites, and enhance operational control. Such configurations also improve scalability and resilience, as they allow new suppliers to be integrated through local connections to existing hubs and enable rerouting in response to disruptions. In the context of global automotive manufacturing, this structure supports efficient coordination across multiple tiers, regions, and transport modes. However, fully realizing these benefits requires advanced transportation planning capabilities to manage the resulting complexity in routing, consolidation, and timing decisions.
@@
About 80% of them connect suppliers to industrial sites and typically last less than two weeks. Another 15%, mostly intra-continental, either originate from or terminate at platforms. The remaining legs connect platforms to each other and account for the rest of the intra-continental links and nearly all inter-continental ones, with durations ranging from three to eight weeks. These platform–platform legs often rely on long-haul modes such as deep-sea shipping and require early capacity reservation and schedule coordination. The heterogeneity in leg types has important operational implications. Supplier–site legs are typically managed within short-term tactical planning horizons, whereas platform–to-platform connections involve strategic decisions regarding routing, frequency, and transport capacity. Moreover, intercontinental legs–though less frequent–represent a disproportionately large share of transportation cost and carbon emissions. Consolidating flows at platforms is therefore essential not only for cost efficiency but also for meeting sustainability targets.
In summary, the network features a distinctive topology: a compact, high-density core that enables efficient long-haul transport and consolidation, and a vast, low-connectivity periphery that generates most of the supply variability. These structural characteristics impose specific methodological requirements on transportation planning models and motivate the modeling approach developed in the following sections.

### v3

@@ Abstract
We study the shipper-side design of large-scale inbound transportation networks, motivated by the global supply chain of the carmaker Renault. We formalize the Shipper Transportation Planning Problem (STPP), which integrates discrete flow consolidation via explicit bin-packing, time-expanded routing, and operational regularity constraints. To solve this high-complexity combinatorial problem at an industrial scale, we propose a tailored Iterated Local Search (ILS) metaheuristic. The algorithm combines large-neighborhood search with MILP-based perturbations and leverages bundle-specific decompositions to obtain scalable lower bounds and effective search guidance. Computational experiments on real industrial data involving more than 700,000 commodities and 1.2 million arcs demonstrate that the ILS achieves an average gap of 7.9% to the best available lower bound. The results reveal a 23.2% cost-reduction potential compared to legacy planning benchmarks. Most significantly, the proposed framework is currently deployed in production at Renault, where it supports weekly strategic decisions and generates realized cost savings estimated at approximately €20 million per year. Our analysis yields key managerial insights: we demonstrate that explicit 1D bin-packing is a critical step forward for realistic consolidation modeling, that transport regularity offers a robust balance between cost and stability, and that high-volume global networks benefit significantly from in-house strategic planning over third-party outsourcing. To the best of our knowledge, this is the first work to successfully solve a shipper-side transportation design problem at this magnitude.
@@ 1 Introduction
Renault exemplifies such a case. As a global car manufacturer with a vast supplier base and worldwide production footprint, Renault moves millions of parts across continents each year. This results in an inbound logistics network of exceptional size and complexity—comprising thousands of suppliers, hundreds of logistics platforms, and multiple industrial sites—requiring highly detailed and scalable optimization methods to ensure efficient operations.
This paper arises from a close industrial collaboration with Renault and addresses a real-world strategic problem: the long-term planning of the company’s inbound transportation network. In this context, Renault initiated a strategic effort to modernize and strengthen its inbound logistics planning capabilities. Recent initiatives have focused on redesigning planning processes around systematic use of operational data and advanced decision‑science tools to enhance transparency and network performance. This initiative has led to the development of multiple tools adapted to each planning level: strategic, tactical and operational. This project was launched after the company’s internal strategic transportation planning engine proved unable to deliver satisfactory results on worldwide‑scale instances, particularly due to the complexity of consolidation mechanisms and the scale of the network. As a result, Renault required a new generation of planning tools—beyond off‑the‑shelf software or generic outsourcing models—built on bespoke, high‑performance optimization methods capable of handling the full global problem, capturing the system’s specific structural characteristics, incorporating finer consolidation models, and producing high‑quality solutions within reasonable computation times.
In this work, our objective is to produce a provisional transportation plan for parts to be assembled in the next six months. Each commodity—defined as a part to be delivered from a specific supplier to a specific industrial site by a given date—must be routed through Renault’s logistics network in a cost-efficient manner. The decision involves choosing whether to ship parts directly or to consolidate them through intermediate platforms. To allow for optimized decision-making in this context, we formalize the underlying Shipper Transportation Planning Problem (STPP), tailored to the needs of a high-volume industrial shipper, and develop a scalable algorithmic framework, capable of generating high-quality plans on real data from one of Europe’s largest manufacturers.
@@ 1.1 State of the art
The design and optimization of transportation networks have been long-standing challenges in operations research, especially in the context of logistics and supply chain management. The foundational modeling framework is that of multicommodity network flow problems, extensively studied since the 1960s (ahuja1988network). Early research already recognized the critical importance of flow consolidation in freight transportation (powell1983load). The service network design (SND) literature, a broad class of planning problems that captures consolidation-based freight operations, is the appropriate reference point for our problem. It was formally established in crainic2000service and further analyzed by andersen2009service and crainic2020exact. More recently, crainicNetworkDesignApplications2021a crainic202550 provide comprehensive surveys of the methodological landscape.
Within this family of problems, our work extends a scheduled multi-layer service network design problem with unsplittable flows by incorporating explicit bin-packing consolidation, flexible delivery times, additional regularity constraints, and a strategic shipper-oriented viewpoint, all within an extremely large-scale industrial setting. Among the various SND variants, we focus primarily on load plan design problems (LPDPs), which share many structural similarities with our formulation while remaining more tractable than full-fledged SND models.
Table 1 compares our model with representative works in the SND and LPDP literature: [1] jarrah2009, [2] lindsey2016improved, [3] eomrecursive, [4] he2022exact and [5] zhu2014scheduled.
Table 1: Similarities and differences with the most relevant related work.
Characteristics [1] [2] [3] [4] [5] Our work
Method Heuristic Heuristic Heuristic Exact Heuristic Heuristic
Flow type U-MCF U-MCF U-MCF U-MCF U-MCF U-MCF
Unrestricted path set × × ×
Network nodes (time-space) 5100 5400 1000 4300 210 63 500
Network arcs (time-space) 210 000 864 000 248 000 135 000 6100 1 190 000
Number of commodities 680 60 000 40 000 1000 1000 700 000
Time horizon 1 week 4 days 1 day 3 hours 1 week 6 months
Bin-Packing consolidation × × × × $\sim$
Regularity type T-G-IT T-G-IT × × G T-G
Resource management × × × ×
Flexible delivery time × × × × ×
Point of view Carrier Carrier Carrier Carrier Carrier Shipper
Decision scope Tactic Tactic Tactic Operations Tactical Strategic
Lower Bound × ×
Notes Usage: U = Unsplittable, MCF = Multi-Commodity Flow, T = Time regularity, G = Group regularity, IT = In-Tree regularity.
@@
Exact solution methods for such problems typically rely on mathematical programming techniques, including decomposition methods such as Lagrangian relaxation, Benders decomposition, and Dantzig-Wolfe reformulation (crainic2020exact) (boland2017continuous; fontaine2021scheduled; he2022exact) [4]. These methods are effective on moderate-size instances but struggle to scale. Studies by gendron2014branch and frangioni2017computational report computation times of several hours even on modest networks, involving a few hundred arcs and commodities.
To address larger systems, the literature increasingly relies on heuristic and matheuristic techniques. Metaheuristics such as large neighborhood search or population-based approaches have demonstrated strong empirical performance on realistically sized problems (crainic2021heuristics; kazemzadeh2022node; gendron2018matheuristics; paraskevopoulos2016cycle) (jarrah2009) [1]. In road transportation planning, for instance, erera2013improved and lindsey2016improved [2] solve instances involving tens of thousands of arcs and commodities by combining local search strategies with sophisticated descent-based matheuristics. More recently, eomrecursive [3] proposed a recursive partitioning and batching strategy to improve scalability through problem decomposition. Our present work, with massive commodity volumes, demands algorithmic solutions that scale even further.
Explicit consolidation decisions have received far less attention. To the best of our knowledge, no LPDP formulation integrates bin-packing constraints to determine the number of services required on each arc (bakir2021motor). A partial exception occurs in train blocking and scheduling models (zhu2014scheduled; kienzle2025intermodal), where commodities are first grouped into blocks, which are then assigned to trains. While this two-stage consolidation mimics bin-packing, block structures are precomputed and therefore eliminate flexibility in commodity-to-block assignments. The absence of explicit bin-packing constraints is noteworthy: introducing them dramatically increases problem size and complicates decomposition and matheuristic strategies. By explicitly modeling discrete 1D bin-packing, we move beyond the standard continuous volume relaxations (aggregate capacity bounds) prevalent in the literature, capturing the combinatorial necessity of filling discrete transport units—a primary cost driver in industrial logistics.
Finally, existing LPDP models overwhelmingly adopt a carrier-side viewpoint, embedding resource management constraints and fixed departure schedules (lindsey2016improved). In contrast, a shipper-oriented setting removes resource constraints, since insufficient capacity from one carrier may be compensated by another, and permits relaxed scheduling assumptions. This shift loosens regularity requirements that dramatically expands that search space of such problems and breaks methods based on enumerating feasible paths.
Overall, this review highlights a significant gap: while LPDPs and SND problems have been widely studied from the perspective of carriers and third‑party logistics providers, the shipper’s perspective—combined with explicit consolidation decisions and massive commodity volumes for which current approaches remain insufficient—remains largely unexplored.
@@ 1.2 Contributions
This paper studies the shipper-side design of large-scale inbound transportation networks, focusing on the strategic planning problem faced by Renault. Building on the gaps identified in the service network design (SND) and load plan design (LPDP) literature, we make four main contributions:
(i) A novel shipper problem setting within the SND family.
We formalize the Shipper Transportation Planning Problem (STPP), which integrates several structural features rarely combined in prior studies: explicit bin-packing consolidation, flexible delivery-time windows, regularity constraints, and massive-scale unsplittable flows. Unlike carrier-centric SND models, our formulation adopts a strategic shipper-oriented viewpoint that removes fixed fleet resource constraints while introducing complex requirements for operational stability and long-haul synchronization. To our knowledge, this is the first formulation in the literature that simultaneously incorporates all these elements at an industrial scale.
(ii) A scalable metaheuristic framework tailored to this setting.
Methodologically, we develop a novel Iterated Local Search (ILS) metaheuristic specifically designed to navigate the STPP’s combinatorial complexity. The framework integrates: (a) specialized large-neighborhood operators adapted to bin-packing–based consolidation, (b) perturbation mechanisms guided by tractable MILP relaxations, and (c) bundle-based decompositions that yield scalable lower bounds. This approach enables the discovery of high-quality solutions for instances involving 700,000 commodities and 1,200,000 arcs–a scale where standard SND heuristics and exact decomposition methods, such as Column Generation, reach their practical limits.
(iii) A computational study at industrial scale with realized impact.
Empirically, we demonstrate that the proposed ILS achieves substantial performance improvements over both Renault’s legacy planning solutions and state-of-the-art benchmark heuristics. On the largest “World” instances, the algorithm delivers solutions revealing a 23.2% cost-reduction potential relative to previous operational plans. Importantly, the solution has been deployed in production at Renault, where it currently supports weekly planning decisions and generates realized cost savings estimated at approximately €20 million per year. To support reproducibility, we provide an anonymized version of the industrial data used in this study.
(iv) Managerial insights for global inbound logistics.
Our analysis provides actionable guidance for large-scale industrial shippers: (a) Consolidation Fidelity: We prove that explicit 1D bin-packing is a “step forward” for realistic planning; relying on standard “giant container” relaxations leads to cost distortions of approximately 20%. (b) Price of Regularity: We quantify the trade-off between operational stability and cost, showing that highly regular plans align with practitioner preferences while sacrificing only marginal potential savings. (c) Strategic Insourcing: We identify a volume-dependent threshold for outsourcing, demonstrating that for high-volume global networks, in-house planning consistently outperforms 3PL models.
Overall, our work extends the scope of the service network design literature by introducing a shipper-side, bin-packing–explicit formulation and providing the first computational evidence that such problems can be solved effectively at a global industrial scale.
@@ 2 Case study & Problem Setting
This section presents the empirical and operational context for the transportation planning problem studied in this paper. We focus on Renault’s inbound supply network—a global, high-volume, multi-tier logistics system. The section characterizes the network’s topology, flow structure, and planning requirements, beginning with its physical architecture and consolidation logic, followed by key flow characteristics and logistical challenges. A detailed and extended version of this case study is available in Appendix A. Finally, we present an informal problem statement that synthesizes these elements into a shipper-side transportation planning problem, capturing the scale, constraints, and cost trade-offs specific to this context.
@@ 2.1 Supply Network
These maps highlight the global scale and complexity of Renault’s supply network, which comprises more than 3,000 sites, 40 industrial sites, and 100 logistics platforms. Most sites are located in Europe, and suppliers account for the vast majority of nodes. This scale, combined with heterogeneous lead times—ranging from a few days for regional flows to several weeks for intercontinental shipments—creates significant planning complexity, amplified by exposure to disruptions and the need to coordinate consolidation across multiple tiers. The network follows a hub‑and‑spoke structure: suppliers form a sparsely connected periphery, while platforms act as high‑degree hubs enabling consolidation and reducing the number of direct supplier–plant connections. Such configurations support economies of scale while also improve scalability and resilience, as they allow new suppliers to be integrated through local connections to existing hubs and enable rerouting in response to disruptions. However, fully realizing these benefits requires sophisticated planning tools to manage consolidation, timing, and routing decisions effectively. Network patterns further reflect this structure. Most legs connect suppliers to industrial sites and have short durations; platform‑to‑platform legs, although fewer, are long‑haul, strategically planned, and account for a disproportionate share of transportation cost and emissions. Consolidation at platforms is therefore essential for cost efficiency and sustainability, especially for intercontinental flows. In this work, each leg is served by a single mode (trucks inland, vessels overseas), reflecting data availability constraints, though the model is fully compatible with a multimodal settings-as discussed in Appendix B. Overall, the network combines a high‑density consolidation core with a vast, low‑connectivity supplier periphery, a topology that drives the methodological requirements motivating the modeling framework developed in the following sections.